I need help solving this fractional equation:
5k/48 - 2k/9 = 17/18 (these are suppose to be fractions but not sure how to get the line under the top numbers)
thank you in advance. I really appreciate.

Answer by Alan3354(69443)

I need help solving this fractional equation:
5k/48 - 2k/9 = 17/18
----------------------
Step 1, get rid of fractions.
Multiply thru by 144/144
15k - 32k = 136
-17k = 136
k = -8
--------